Today, I started to check the transfer functions of the tower part.
I measured the IP and Yaw_BF TFs using the Fujii's templates located in: users/VISsvn/TypeAtower/ITMY/measuremnt/TF/template.
Since the oscillation of the IP and BF was to small, I incresed of 3 times the noise injected respect to the default one:
- IP: 3000 counts insted of 1000 counts for L and T and 1500 counts insted of 500 counts for Yaw.
- BF_Y: 3000 counts instead of 800 counts.
During all measuremens the optical levers signals were in the range.
Pic 1, Pic2, Pic3 show the IP transfer functions: today TFs (red curves) vs June 2019: The IP is healthy.
Pic 4 shows the BF_Y TF : Today TF (red curve) vs Febrary 2019 (blue curve): BF_Y body looks not touching and not rubbing (healthy).
I also measured the GAS TFs and compared with the old ones measured in March 2019 (yellow curve).
- GAS filters: F1 (pic5), F2 (pic6) , F3 (pic7), BF (pic8)
The GAS Filters are moving, but the position of the main peacks as well as the shape and the DC gain are different (Pic8, Pic6).
When I turned on the IP loops I observed that the IP_Y loop was instable, so I changed the gain and thye DC control: now is stable.
I turned on all IP and BF_Y loops.
Concernig the GAS loops I did't turned on them, because of the new TFs shape I need time to reshape the control filters.
Now the health check of the Paylod part is ongoing.
The .xml file of the measurements are stored in: